# Hierarchical Adaptive networks with Task vectors for Test-Time Adaptation

### v1

@@ Abstract
Test-time adaptation allows pretrained models to adjust to incoming data streams, addressing distribution shifts between source and target domains. However, standard methods rely on single-dimensional linear classification layers, which often fail to handle diverse and complex shifts. We propose Hierarchical Adaptive Networks with Task Vectors (Hi-Vec), which leverages multiple layers of increasing size for dynamic test-time adaptation. By decomposing the encoder’s representation space into such hierarchically organized layers, Hi-Vec, in a plug-and-play manner, allows existing methods to adapt to shifts of varying complexity. Our contributions are threefold: First, we propose dynamic layer selection for automatic identification of the optimal layer for adaptation to each test batch. Second, we propose a mechanism that merges weights from the dynamic layer to other layers, ensuring all layers receive target information. Third, we propose linear layer agreement that acts as a gating function, preventing erroneous fine-tuning by adaptation on noisy batches. We rigorously evaluate the performance of Hi-Vec in challenging scenarios and on multiple target datasets, proving its strong capability to advance state-of-the-art methods. Our results show that Hi-Vec improves robustness, addresses uncertainty, and handles limited batch sizes and increased outlier rates.
@@ 1 Introduction
The test-time adaptation paradigm was introduced to mitigate performance degradation of deep learning models when faced with new data that differ from the training distribution [83, 39, 57, 86, 53]. These techniques adjusts the model’s parameters or internal representations on test data, aiming to handle the distribution gap between training and inference, and to manage previously unseen shifts. Recent advances show that test-time adaptation can effectively tackle challenges such as unexpected image corruptions [83, 53, 11] and domain changes [89, 39, 6], in tasks like classification [57, 48] and semantic segmentation [7, 80].
A major challenge in test-time adaptation originates from the limited flexibility of representations learned by standard deep learning methods, which typically use a single linear layer after the encoder to project these representations into the output space, as shown in Fig. 1a. This setup forces all features through a single projection, leading to information diffusion [77], where fine-grained details become blurred across layers. Additionally, the final layer of the encoder may experience rank collapse, a phenomenon in which representations lose expressivity and become overly constrained [23, 47].
In response to the aforementioned drawbacks, Kusupati et al. [47] propose Matryoshka Representation Learning (MRL), which attaches multiple linear layers of different sizes to the encoder. Each layer in MRL maps a progressively larger subset of the encoder’s features to the output space. This hierarchical design promotes coarse-to-fine feature extraction, enhancing the overall expressivity and flexibility of the learned representation.
Building on MRL, we introduce Hi-Vec (Hierarchical Layers with Task Vectors), a novel framework designed to handle diverse distribution shifts during test time. Hi-Vec leverages hierarchical representations to dynamically identify the optimal level of granularity required to handle the current distribution shift (see also Fig. 1b). This forms a versatile backbone for test-time adaptation, accommodating a broad range of data distributions.
@@
In Section 3, we introduce the problem setup and hierarchical linear layer configuration. Section 4 presents our method’s theoretical foundation and its three main components. Section 5 details experiments integrating Hi-Vec with four state-of-the-art test-time adaptation methods. We evaluate two challenging scenarios: (i) Section 5.1 assesses outlier-aware settings, where test batches mix target samples with out-of-distribution ones from unrelated datasets, simulating real-world inference with unknown classes or domains. (ii) Section 5.2 examines spurious correlations, using target datasets with non-causal links (e.g., background features tied to classes) differing from the source domain, forcing adaptation without misleading source features. In Section 5.3, we detail the benefits and insights of using our framework.
@@ 2 Related Work
Test-Time Adaptation. Test-time adaptation allows models to address distribution shifts by adapting and predicting simultaneously on incoming streams of unlabeled target data [10, 12, 59, 90, 74, 88, 54, 92, 91, 9, 42, 81, 5, 96, 50, 21, 2]. Unlike domain adaptation [82, 36] and domain generalization [60, 29], test-time adaptation uniquely operates during inference, making it suitable for real-world scenarios where target data is unavailable during training and arrives sequentially. Standard fine-tuning approaches typically adjust a fixed set of parameters, such as batch normalization statistics [100, 61], full model parameters [83, 53], or single-dimensional linear layers [41, 39, 102]. Beyond these, recent works have explored overall layer selection [74] and surgical fine-tuning [49] for targeted adaptation. These methods have been evaluated on unseen scenarios, including noise-corrupted datasets [83, 27] and datasets with varying domain information [39, 53]. Recently, Yu et al. [97] incorporated outliers into the evaluation protocol and introduced a memory-based approach with self-weighted entropy to handle diverse distribution shifts. However, like standard test-time adaptation methods, their approach relies on a single-dimensional linear layer and repeatedly uses the same batch norm parameters, limiting feature expressivity.
@@ 3.2 Hierarchical Layers
To learn coarse-to-fine representations, we adopt Matryoshka layers [47], which we refer to as hierarchical linear layers. These layers decompose the encoder’s latent representation $\mathbf{z}\in\mathbb{R}^{d}$ into a set of subspaces $\mathbf{z}_{\mathcal{M}}=\{\mathbf{z}_{1:k}\in\mathbb{R}^{k}\mid k\leq d\}$, organizing the feature space into increasingly refined levels. This hierarchical organization enhances adaptability and encourages multi-scale representation learning.
@@ 4 Methodology: Hierarchical Layers for Test-Time Adaptation
Hi-Vec leverages the hierarchical coarse-to-fine structure learned during source training to address diverse distribution shifts in the target domain $\mathcal{T}={D_{t}}$, where different dimensional embeddings capture varying levels of abstraction [56, 47]. Therefore, each hierarchical linear layer $\bm{\phi}_{1},\bm{\phi}_{2},\dots,\bm{\phi}_{k}$ encodes a distinct level of representational granularity, making it inherently suitable for adaptive feature learning.
@@ 4.1 Dynamic Layer Selection
To identify the most suitable layer for each batch dynamically, we compute the gradient norm for each layer independently, using an entropy minimization loss [83]. The layer with the lowest gradient norm is selected for prediction and adaptation, as it indicates best alignment. Formally, let $\|\nabla_{W_{\bm{\phi}}}\mathcal{L}\|$ denote the gradient norm with respect to the parameters of layer $\bm{\phi}$. The optimal layer $\bm{\phi}^{*}$ is selected as:
@@
Once $\bm{\phi}^{*}$ is selected, the encoder $\bm{\theta}$ and the chosen hierarchical layer $\bm{\phi}^{*}$ can be fine-tuned using existing test-time adaptation methods [83, 48, 97, 62]. In particular, any test-time method relying on an unsupervised loss for parameter updates is seamlessly compatible with our framework. As a plug-and-play extension, Hi-Vec enhances existing methods through hierarchical layers and dynamic selection promoting targeted adaptation per batch.
@@ 4.2 Target Information sharing with Task Vectors
To allow for dissemination of target information from the selected layer $\bm{\phi}^{*}$ to other layers, we propose a task vectors [37] based mechanism that does not require gradient computation or backpropagation.
Unlike uniform merging of prior works, our framework handles hierarchical layers of varying dimensions through dimension-aware updates, merging subsets of compatible weights. In this framework, each hierarchical layer $\bm{\phi}$ acts as a task vector $\mathbf{v}_{\bm{\phi}}$. We compute affinities using cosine similarity [32] and identify similar layers exceeding a threshold $\tau$:
@@
where $W_{\bm{\phi}}[\bm{\phi}^{*}]$ represents the subset of weights in $W_{\bm{\phi}}$ corresponding to those in $W_{\bm{\phi}^{*}}$, and $\alpha$ is a scaling factor. All other weights in $W_{\bm{\phi}}$ remain unchanged.
This direct propagation of updates from $\bm{\phi}^{*}$ to similar layers is theoretically supported by the principles of linear mode connectivity [25], because the layers are trained jointly on source data, creating smooth, connected paths through low-error regions in the loss landscape, which allows reliable model weight merging.
@@ 4.3 Hierarchical Linear Layer Agreement
We propose a hierarchical linear layer agreement mechanism to reliably detect out-of-distribution (OOD) samples and mitigate model misspecification from noisy data. For detected OOD samples, adaptation (backpropagation) is skipped, and the selected layer $\bm{\phi}^{*}$ is used solely for inference. In contrast, in-distribution (ID) target batches undergo both inference and adaptation (see also Fig. 2).
@@ 5 Experiments and Results
Datasets and shifts: Five target datasets and six outlier datasets. The respective source models are trained on public datasets, i.e. Cifar-10 [45], Cifar-100 [45], Waterbirds [73] and ImageNet [19]. Evaluation is performed on five target datasets and six outlier datasets with both kinds of diverse shifts as provided in  [97, 48, 35]. In the outlier-aware scenario, test batches combine target data from respective corrupted benchmarks Cifar-10-C [35], Cifar-100-C [35], or ImageNet-C [35] with outlier samples from LSUN-C [97], SVHN-C [97], TinyImageNet-C [35], Places-365-C [97], and Textures-C [97]. For spurious correlation scenarios, we use domain-specific datasets such as Waterbirds [73] and ColoredMNIST, with non-causal correlations. We provide additional dataset details in the Supplement.
Implementation Details. We train source models on Cifar-10, Cifar-100, Waterbirds, and ImageNet, respectively, using ResNet-18 and ResNet-50 as encoders. Hierarchical linear layers are integrated with each encoder, mapping features into subspaces of increasing dimensionality starting from 8 up to the encoder’s output dimension (i.e., $m_{i}=2^{i+2}$[47]; 512 for ResNet-18 or 2048 for ResNet-50). Training these layers is straightforward and requires no additional modifications. We evaluate Hi-Vec using the same metrics from [97, 48] and adopting their reported baseline performances. For ablations and additional experiments, we integrate Hi-Vec with re-implemented baselines from their official GitHub repositories. For Hi-Vec, we compute gradient norms with PyTorch’s built-in functions and apply its padding utilities to equalize model weights for similarity calculations. For test-time online adaptation, we adhere to the standard protocol by incrementing the target data consisting of outliers or spurious correlations iteratively and keep updating and evaluating the model following prior works [97, 62, 83, 48]. In the tables, we additionally report vanilla MRL [47] as a zero-shot baseline akin to source model performance, selecting the dimension with the highest accuracy for brevity. Hyperparameters and further details are provided in the Supplement. We will release the code publicly.
@@ 5.1 Results for Adaptation with Outlier Datasets
Table 1 highlights the constraints of baseline methods, including Tent, Sar, and Stamp, on Cifar-10-C and Cifar-100-C with outlier datasets. Hi-Vec’s integration addresses this due to multi-scale linear layers that promote enhanced feature expressivity and dynamic layer selection for adaptive fine-tuning, yielding consistent gains across all metrics. On Cifar-10-C datasets with the four outliers, Hi-Vec’s integrations improve the mean accuracy for all methods by 2.47%. For instance, Hi-Vec + Stamp delivers the highest improvements of +5.7% ACC (to 83.6%), +8.2% AUC (to 91.4%), and +7.2% H-Score (to 87.3%) on Cifar-10-C with Noise; +3.4% ACC (to 85.7%) with SVHN-C; and +3.9% ACC (to 86.5%) with TinyImageNet-C. Similar improvements hold for Cifar-100-C, where Hi-Vec integrations mostly emerge as top performers.
Next, we also provide results on large-scale datasets, showcasing Hi-Vec’s effectiveness on ImageNet using Textures and Places-365 as outlier target datasets. Table 3 presents results across 15 different image corruptions as in  [35]. Leveraging its adaptive mechanisms, Hi-Vec enables baseline approaches to advance their performance even further, thereby addressing large-scale outlier datasets.
@@ 6 Conclusion
We proposed Hi-Vec, a novel framework that integrates with state-of-the-art test-time adaptation methods to allow them to address diverse shifts at test-time. By leveraging hierarchical linear layers, Hi-Vec learns coarse-to-fine representations that can be seamlessly integrated with existing methods. We further advance the architecture by integrating dynamic layer selection for precise adaptation, target information sharing for cross-layer coherence, and layer agreement to detect noisy samples and skip adaptation on noisy test batches. We have evaluated our framework, Hi-Vec, in challenging scenarios, including outlier shifts and spurious correlations. Our results demonstrate the strong capabilities of Hi-Vec, consistently enhancing existing methods across diverse shifts. Additionally, our framework significantly improves baseline methods in handling small batch sizes, addressing high outlier samples in target data, mitigating uncertainty, and supporting both standard and open-set test-time adaptation settings, demonstrating its versatility for real-world challenges. The strong performance of Hi-Vec motivates us to further enhance its capabilities. Currently, Hi-Vec relies on backpropagation to select the optimal hierarchical layer, limiting its applicability to optimization-based methods. For future work, we also aim to address parameter-free adaptation methods, thereby broadening the range of test-time adaptation methods that can be integrated.

### v2

@@ Abstract
Test-time adaptation allows pretrained models to adjust to incoming data streams, addressing distribution shifts between source and target domains. However, standard methods rely on single-dimensional linear classification layers, which often fail to handle diverse and complex shifts. We propose Hierarchical Adaptive Networks with Task Vectors (Hi-Vec), which leverages multiple layers of increasing size for dynamic test-time adaptation. By decomposing the encoder’s representation space into such hierarchically organized layers, Hi-Vec, in a plug-and-play manner, allows existing methods to adapt to shifts of varying complexity. Our contributions are threefold: First, we propose dynamic layer selection for automatic identification of the optimal layer for adaptation to each test batch. Second, we propose a mechanism that merges weights from the dynamic layer to other layers, ensuring all layers receive target information. Third, we propose linear layer agreement that acts as a gating function, preventing erroneous fine-tuning by adaptation on noisy batches. We rigorously evaluate the performance of Hi-Vec in challenging scenarios and on multiple target datasets, proving its strong capability to advance state-of-the-art methods. Our results show that Hi-Vec improves robustness, addresses uncertainty, and handles limited batch sizes and increased outlier rates. Code: https://github.com/ambekarsameer96/Hi-Vec
@@ 1 Introduction
The test-time adaptation paradigm was introduced to mitigate performance degradation of deep learning models when faced with new data that differ from the training distribution [82, 38, 56, 85, 52]. These techniques adjusts the model’s parameters or internal representations on test data, aiming to handle the distribution gap between training and inference, and to manage previously unseen shifts. Recent advances show that test-time adaptation can effectively tackle challenges such as unexpected image corruptions [82, 52, 9] and domain changes [88, 38, 4, 2], in tasks like classification [56, 47] and semantic segmentation [5, 79].
A major challenge in test-time adaptation originates from the limited flexibility of representations learned by standard deep learning methods, which typically use a single linear layer after the encoder to project these representations into the output space, as shown in Fig. 1a. This setup forces all features through a single projection, leading to information diffusion [76], where fine-grained details become blurred across layers. Additionally, the final layer of the encoder may experience rank collapse, a phenomenon in which representations lose expressivity and become overly constrained [22, 46]. In response to the aforementioned drawbacks, Kusupati et al. [46] propose Matryoshka Representation Learning (MRL), which attaches multiple linear layers of different sizes to the encoder. Each layer in MRL maps a progressively larger subset of the encoder’s features to the output space. This hierarchical design promotes coarse-to-fine feature extraction, enhancing the overall expressivity and flexibility of the learned representation.
Building on MRL, we introduce Hi-Vec (Hierarchical Layers with Task Vectors), a novel framework designed to handle diverse distribution shifts during test time. Hi-Vec leverages hierarchical representations to dynamically identify the optimal level of granularity required to handle the current distribution shift (see also Fig. 1b). This forms a versatile backbone for test-time adaptation, accommodating a broad range of data distributions. Our key contributions are as follows:
@@
In Section 3, we introduce the problem setup and hierarchical linear layer configuration. Section 4 presents our method’s theoretical foundation and its three main components. Section 5 details experiments integrating Hi-Vec with four state-of-the-art test-time adaptation methods. We evaluate two challenging scenarios: (i) Section 5.1 assesses outlier-aware settings, where test batches mix target samples with out-of-distribution ones from unrelated datasets (ii) Section 5.2 examines spurious correlations, using target datasets differing from the source domain, forcing adaptation without misleading on source features. In Section 5.3, we detail the benefits and insights of using our framework.
@@ 2 Related Work
Test-Time Adaptation. Test-time adaptation allows models to address distribution shifts by adapting and predicting simultaneously on incoming streams of unlabeled target data [8, 10, 58, 89, 73, 87, 53, 91, 90, 7, 41, 80, 3, 95, 49, 20]. Unlike domain adaptation [67, 81, 35] and domain generalization [59, 28], test-time adaptation uniquely operates during inference, making it suitable for real-world scenarios where target data is unavailable during training and arrives sequentially. Standard fine-tuning approaches typically adjust a fixed set of parameters, such as batch normalization statistics [99, 60], full model parameters [82, 52], or single-dimensional linear layers [40, 38, 101]. Beyond these, recent works have explored overall layer selection [73] and surgical fine-tuning [48] for targeted adaptation. These methods have been evaluated on unseen scenarios, including noise-corrupted datasets [82, 26] and datasets with varying domain information [38, 52]. Recently, Yu et al. [96] incorporated outliers into the evaluation protocol and introduced a memory-based approach with self-weighted entropy to handle diverse distribution shifts. However, like standard test-time adaptation methods, their approach relies on a single-dimensional linear layer and repeatedly uses the same batch norm parameters, limiting feature expressivity.
@@ 3.2 Hierarchical Layers
To learn coarse-to-fine representations, we adopt Matryoshka layers [46], which we refer to as hierarchical linear layers due to their increasing dimensionality and representational design. These layers decompose the encoder’s latent representation $\mathbf{z}\in\mathbb{R}^{d}$ into a set of subspaces $\mathbf{z}_{\mathcal{M}}=\{\mathbf{z}_{1:k}\in\mathbb{R}^{k}\mid k\leq d\}$, organizing the feature space into increasingly refined levels. This hierarchical organization enhances adaptability and encourages multi-scale representation learning.
@@ 4 Methodology: Hierarchical Layers for Test-Time Adaptation
Hi-Vec leverages the hierarchical coarse-to-fine structure learned during source training to address diverse distribution shifts in the target domain $\mathcal{T}={D_{t}}$, where different dimensional embeddings capture varying levels of abstraction [55, 46]. Therefore, each hierarchical linear layer $\bm{\phi}_{1},\bm{\phi}_{2},\dots,\bm{\phi}_{k}$ encodes a distinct level of representational granularity, making it inherently suitable for adaptive feature learning [46].
@@ 4.1 Dynamic Layer Selection
To identify the most suitable layer for each batch dynamically, we compute the gradient norm for each layer independently, using an unsupervised entropy minimization loss [82]. The layer with the lowest gradient norm is selected for prediction and adaptation, as it indicates best alignment. Formally, let $\|\nabla_{W_{\bm{\phi}}}\mathcal{L}\|$ denote the gradient norm with respect to the parameters of layer $\bm{\phi}$. The optimal layer $\bm{\phi}^{*}$ is selected as:
@@
Once $\bm{\phi}^{*}$ is selected, the encoder $\bm{\theta}$ and the chosen hierarchical layer $\bm{\phi}^{*}$ can be fine-tuned using existing test-time adaptation methods [82, 47, 96, 61]. In particular, any test-time method relying on an unsupervised loss for parameter updates is seamlessly compatible with our framework. As a plug-and-play extension, Hi-Vec enhances existing methods through hierarchical layers and dynamic selection promoting targeted adaptation per batch. We detail the rationale for choosing the lower gradient norm, since it identifies the layer requiring minimal adaptation in Supplementary A.4.
@@ 4.2 Target Information sharing with Task Vectors
To allow for dissemination of target information from the selected layer $\bm{\phi}^{*}$ to other layers, we propose a task vectors [36] based mechanism that does not require gradient computation or backpropagation. Unlike uniform merging of prior works, our framework handles hierarchical layers of varying dimensions through dimension-aware updates, merging subsets of compatible weights. In this framework, each hierarchical layer $\bm{\phi}$ acts as a task vector $\mathbf{v}_{\bm{\phi}}$. We compute affinities using cosine similarity [31] and identify similar layers exceeding a threshold $\tau$:
@@
where $W_{\bm{\phi}}[\bm{\phi}^{*}]$ represents the subset of weights in $W_{\bm{\phi}}$ corresponding to those in $W_{\bm{\phi}^{*}}$, and $\alpha$ is a scaling factor. All other weights in $W_{\bm{\phi}}$ remain unchanged. A detailed formulation of the notation introduced here can be found in Supplementary A.9. This direct propagation of updates from $\bm{\phi}^{*}$ to similar layers is theoretically supported by the principles of linear mode connectivity [24], because the layers are trained jointly on source data, creating smooth, connected paths through low-error regions in the loss landscape, which allows reliable model weight merging.
@@ 4.3 Hierarchical Linear Layer Agreement
We propose a hierarchical linear layer agreement mechanism to reliably detect out-of-distribution (OOD) samples and mitigate model misspecification from noisy data. For detected OOD samples, adaptation (backpropagation) is skipped, and the selected layer $\bm{\phi}^{*}$ is used solely for inference. In contrast, in-distribution (ID) target batches undergo both inference and adaptation (see also Fig. 2). Notably, skipping of adaptation on OOD samples aligns with the procedure established in EATA [61] and SAR [62], which have been proven that adapting to OOD data harms performance during test-time.
@@ 5 Experiments and Results
Datasets and shifts: Five target datasets and six outlier datasets. The respective source models are trained on public datasets, i.e. Cifar-10 [44], Cifar-100 [44], Waterbirds [72] and ImageNet [17]. Evaluation is performed with diverse shifts as provided in  [96, 47, 34]. In the outlier-aware scenario, test batches combine target data from respective corrupted benchmarks Cifar-10-C [34], Cifar-100-C [34], or ImageNet-C [34] with outlier samples from LSUN-C [96], SVHN-C [96], TinyImageNet-C [34], Places-365-C [96], and Textures-C [96]. For spurious correlation scenarios, we use domain-specific datasets such as Waterbirds [72] and ColoredMNIST, with non-causal correlations. We provide additional dataset details in the Supplement.
Implementation Details. Hierarchical linear layers are integrated with each encoder, mapping features into subspaces of increasing dimensionality starting from 8 up to the encoder’s output dimension (i.e., $m_{i}=2^{i+2}$[46]; 512 for ResNet-18 or 2048 for ResNet-50). Training these layers is straightforward and requires no additional modifications. We evaluate Hi-Vec using the same metrics from [96, 47] and adopting their reported baseline performances. For Hi-Vec, we compute gradient norms with PyTorch’s built-in functions and apply its padding utilities to equalize model weights for similarity calculations. For test-time online adaptation, we adhere to the standard protocol by incrementing the target data consisting of outliers or spurious correlations iteratively and keep updating and evaluating the model following prior works [96, 61, 82, 47]. The direct integration of MRL [46] linear layers with baseline TTA methods is not feasible due to their multi-head structure. Consequently, in the tables, we additionally report vanilla MRL [46] as a zero-shot baseline akin to source model performance, selecting the dimension with the highest accuracy for brevity. Hyperparameters and further details are provided in the Supplement. We will release the code publicly.
@@ 5.1 Results for Adaptation with Outlier Datasets
Table 1 highlights the constraints of baseline methods, including Tent, Sar, and Stamp, on Cifar-10-C and Cifar-100-C with outlier datasets. Hi-Vec’s integration addresses this due to multi-scale linear layers that promote enhanced feature expressivity and dynamic layer selection for adaptive fine-tuning, yielding consistent gains across all metrics. On Cifar-10-C datasets with the four outliers, Hi-Vec’s integrations improve the mean accuracy for all methods by 2.47%. For instance, Hi-Vec + Stamp delivers the highest improvements of +5.7% ACC (to 83.6%), +8.2% AUC (to 91.4%), and +7.2% H-Score (to 87.3%) on Cifar-10-C with Noise; +3.4% ACC (to 85.7%) with SVHN-C; and +3.9% ACC (to 86.5%) with TinyImageNet-C. Similar improvements hold for Cifar-100-C, where Hi-Vec integrations mostly emerge as top performers. Next, we also provide results on large-scale datasets, showcasing Hi-Vec’s effectiveness on ImageNet using Textures and Places-365 as outlier target datasets. Table 3 presents results across 15 different image corruptions as in  [34]. Leveraging its adaptive mechanisms, Hi-Vec enables baseline approaches to advance their performance even further, thereby addressing large-scale outlier datasets.
@@ 6 Conclusion
We proposed Hi-Vec, a novel framework that integrates with state-of-the-art test-time adaptation methods to allow them to address diverse shifts at test-time. By leveraging hierarchical linear layers, Hi-Vec learns coarse-to-fine representations that can be seamlessly integrated with existing methods. We further advance the architecture by integrating dynamic layer selection for precise adaptation, target information sharing for cross-layer coherence, and layer agreement to detect noisy samples and skip adaptation on noisy test batches. We have evaluated our framework, Hi-Vec, in challenging scenarios, including outlier shifts and spurious correlations. Our results demonstrate the strong capabilities of Hi-Vec, consistently enhancing existing methods across diverse shifts. Additionally, our framework significantly improves baseline methods in handling small batch sizes, addressing high outlier samples in target data, mitigating uncertainty, and supporting both standard and open-set test-time adaptation settings, demonstrating its versatility for real-world challenges. The strong performance of Hi-Vec motivates us to further enhance its capabilities. Notably, eventhough Hi-Vec relies on backpropagation, it can be extended to parameter-free adaptation methods but further research is needed.